OpenStudy (anonymous):

find[1/2(cos 100◦ + isin 100◦)]·[4/5(cos 300◦ + isin 300◦)]and write the result in trigonometric form.

OpenStudy (mertsj):

\[[\frac{1}{2}(\cos 100+i \sin 100)][\frac{4}{5}(\cos 300+i \sin 300)]=\frac{2}{5}(\cos 400+isin400)\]